OpenStudy (anonymous):

Part 1: Using complete sentences, explain which method you would use to solve the following system of equations and why you chose that method. (3 points) 3x + 2y = 7 4x – 3y = –2 Part 2: Show your work and provide the solution to the system. (4 points)

OpenStudy (radar):

I would choose the elimination method because the two original equations do not have x or y with a 1 coefficient. Part 2. Multiply the first equation by 4 as I intend to eliminate the x. 12x + 8y = 28. Now multiply the second equation by 3 getting: 12x - 9y=-6 Now subtract the two adjusted equations ------------ 0x 17y=34 divide by 17 getting y=2 substitute for y in first original equation 3x + 2(2) = 7 3x + 4 =7 3x=7-4 = 3 x=1

OpenStudy (anonymous):

how did you get 17?

OpenStudy (radar):

8y - (-9y) = 17y (subtracting a negative)